# Obstructions for quantitative measure equivalence between locally compact groups

### v1

@@ 6 Absence of quantitatively critical measure equivalence couplings
and we would like to know if this upper bound can be reached. For discrete amenable groups, the first-named author proved that such upper bound cannot be achieved [Cor25, Theorem B], and we want to generalize it to the more general setting of locally compact compactly generated groups.
We need mild assumptions to get the inequality (15). First, we have to assume that $j_{1,H}$ is injective, so as to consider its inverse. We know that the isoperimetric profile is asymptotically equivalent to an injective map (see e.g. [Cor25, Remark 1.2]) so, in some sense, we may assume without loss of generality that the isoperimetric profile is injective. Secondly, the inequality $\varphi\circ j_{1,H}\preccurlyeq j_{1,G}$ means that there exists a constant $C>0$ such that $\varphi\circ j_{1,H}(t)=O\left(j_{1,G}(Ct)\right)$, which means that $\varphi(t)=O\left(j_{1,G}(Cj_{1,H}^{-1}(t))\right)$. To get rid of this constant and obtain the desired composition $j_{1,G}\circ j_{1,H}^{-1}$, we will assume that the isoperimetric profile of $G$ satisfies
@@
This condition already appeared in [Ers03, CD25]. There is no known example of a compactly generated group whose isoperimetric does not satisfy this condition.
Extending the proof of [Cor25, Theorem B] to the case of locally compact compactly generated amenable groups, we want to show Theorem D that we recall here.
@@
Let $G$ and $H$ be non-discrete unimodular locally compact compactly generated groups. Assume that there exist a non-decreasing function $f_{G}$ and an increasing function $f_{H}$ satisfying $f_{G}(n)\simeq j_{1,G}(n)$, $f_{H}(n)\simeq j_{1,H}(n)$ and the following assumptions as $t\to+\infty$:
@@
About Theorem A dealing with volume growth, we prove a similar statement, as in [Cor25, Theorem D]. This is Theorem E that we recall here.
@@
Let $G$ and $H$ be non-discrete unimodular locally compact compactly generated groups. Assume that there exist two increasing functions $f_{G}$ and $f_{H}$ satisfying $f_{G}(n)\simeq V_{G}(n)$, $f_{H}(n)\simeq V_{H}(n)$ and the following assumptions as $t\to+\infty$:
@@
Let $G$ and $H$ be locally compact compactly generated groups, let $G\curvearrowright(X,\nu)$ be a free p.m.p. action on a standard probability space. Let $\varphi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a sublinear and non-decreasing function. If an $H$-valued cocycle $\alpha\colon G\times X\to H$ is $\varphi$-integrable, then there exists a map $\psi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ satisfying the following properties:
@@
For $\varphi\coloneq f_{G}\circ f_{h}^{-1}$, let us assume by contradiction that there exists a $(\varphi,\mathrm{L}^{0})$-integrable measure equivalence coupling from $G$ to $H$. Then Proposition 6.3 implies that this coupling is also $(\psi,\mathrm{L}^{0})$-integrable with a map $\psi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ such that $\psi$ and $t\mapsto t/\psi(t)$ are non-decreasing, and $\psi\not\preccurlyeq f_{G}\circ f_{H}^{-1}$. But Theorem 6.1 implies $\psi\circ j_{1,H}\preccurlyeq j_{1,G}$, namely $\psi\preccurlyeq f_{G}\circ f_{H}^{-1}$ (see the end of the proof of [Cor25, Theorem B] for a justification), so we get a contradiction. ∎
@@
We now proceed to the proof of Proposition 6.3. To this end, we have the following key lemma. In the case of finitely generated groups, this has been proved at the beginning of the proof of [Cor25, Lemma 3.2], which crucially uses finiteness of the generating sets. Here we manage to improve the proof in the non-discrete case.
@@
Let $G$ and $H$ be locally compact compactly generated groups, and let $\varphi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a map. Given a free p.m.p. action $G\curvearrowright(X,\nu)$ on a standard probability space, let $\alpha\colon G\times X\to H$ be a $\varphi$-integrable $H$-valued cocycle. Then there exist a measurable subset $A$ of $G$ and a sequence $(K_{n})_{n\geq 1}$ of positive integers such that:
@@
Let $(K_{n})_{n\geq 1}$ be the sequence provided by Lemma 6.4, together with $A\subset G$ of positive Haar measure as in its statement, and let us denote by $D$ a dense countable set contained in $A$. With the same method as in the proof of [Cor25, Lemma 3.2], we build a map $\psi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ from this sequence, which satisfies the following properties:
@@
Given a compact generating set $S_{G}$ of $G$, it now remains to prove that $\sup_{g\in S_{G}}{\|\alpha(g,.)\|_{\psi}}$ is finite. We use the same techniques as in [BFS13, Appendix A.2 (before Lemma A.1)]. For every $g\in G$, we abusively write

### v2

@@ 6 Absence of quantitatively critical measure equivalence couplings
and we would like to know if this upper bound can be reached. For discrete amenable groups, the first-named author proved that such upper bound cannot be achieved [7, Theorem B], and we want to generalize it to the more general setting of locally compact compactly generated groups.
We need mild assumptions to get the inequality (15). First, we have to assume that $j_{1,H}$ is injective, so as to consider its inverse. We know that the isoperimetric profile is asymptotically equivalent to an injective map (see e.g. [7, Remark 1.2]) so, in some sense, we may assume without loss of generality that the isoperimetric profile is injective. Secondly, the inequality $\varphi\circ j_{1,H}\preccurlyeq j_{1,G}$ means that there exists a constant $C>0$ such that $\varphi\circ j_{1,H}(t)=O\left(j_{1,G}(Ct)\right)$, which means that $\varphi(t)=O\left(j_{1,G}(Cj_{1,H}^{-1}(t))\right)$. To get rid of this constant and obtain the desired composition $j_{1,G}\circ j_{1,H}^{-1}$, we will assume that the isoperimetric profile of $G$ satisfies
@@
This condition already appeared in [12, 6]. There is no known example of a compactly generated group whose isoperimetric does not satisfy this condition.
Extending the proof of [7, Theorem B] to the case of locally compact compactly generated amenable groups, we want to show Theorem D that we recall here.
@@
Let $G$ and $H$ be unimodular locally compact compactly generated groups. Assume that there exist a non-decreasing function $f_{G}$ and an increasing function $f_{H}$ satisfying $f_{G}(n)\simeq j_{1,G}(n)$, $f_{H}(n)\simeq j_{1,H}(n)$ and the following assumptions as $t\to+\infty$:
@@
About Theorem A dealing with volume growth, we prove a similar statement, as in [7, Theorem D]. This is Theorem E that we recall here.
@@
Let $G$ and $H$ be unimodular locally compact compactly generated groups. Assume that there exist two increasing functions $f_{G}$ and $f_{H}$ satisfying $f_{G}(n)\simeq V_{G}(n)$, $f_{H}(n)\simeq V_{H}(n)$ and the following assumptions as $t\to+\infty$:
@@
Let $G$ and $H$ be non-discrete unimodular locally compact compactly generated groups, let $G\curvearrowright(X,\nu)$ be a free measure preserving action on a standard finite-measured space. Let $\varphi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a sublinear and non-decreasing function. If an $H$-valued cocycle $\alpha\colon G\times X\to H$ is $\varphi$-integrable, then there exists a map $\psi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ satisfying the following properties:
@@
By Lemma 2.6, we can assume that the groups are non-discrete. For $\varphi\coloneq f_{G}\circ f_{h}^{-1}$, let us assume by contradiction that there exists a $(\varphi,\mathrm{L}^{0})$-integrable measure equivalence coupling from $G$ to $H$. Then Proposition 6.3 implies that this coupling is also $(\psi,\mathrm{L}^{0})$-integrable with a map $\psi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ such that $\psi$ and $t\mapsto t/\psi(t)$ are non-decreasing, and $\psi\not\preccurlyeq f_{G}\circ f_{H}^{-1}$. But Theorem 6.1 implies $\psi\circ j_{1,H}\preccurlyeq j_{1,G}$, namely $\psi\preccurlyeq f_{G}\circ f_{H}^{-1}$ (see the end of the proof of [7, Theorem B] for a justification), so we get a contradiction. ∎
@@
We now proceed to the proof of Proposition 6.3. To this end, we have the following key lemma. In the case of finitely generated groups, this has been proved at the beginning of the proof of [7, Lemma 3.2], which crucially uses finiteness of the generating sets. Here we manage to improve the proof in the non-discrete case.
@@
Let $G$ and $H$ be non-discrete unimodular locally compact compactly generated groups, and let $\varphi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a map. Given a free measure preserving action $G\curvearrowright(X,\nu)$ on a standard finite measured space, let $\alpha\colon G\times X\to H$ be a $\varphi$-integrable $H$-valued cocycle. Then there exist a measurable subset $A$ of $G$ and a sequence $(K_{n})_{n\geq 1}$ of positive integers such that:
@@
Let $(K_{n})_{n\geq 1}$ be the sequence provided by Lemma 6.4, together with $A\subset G$ of positive Haar measure as in its statement, and let us denote by $D$ a dense countable set contained in $A$. With the same method as in the proof of [7, Lemma 3.2], we build a map $\psi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ from this sequence, which satisfies the following properties:
@@
Given a compact generating set $S_{G}$ of $G$, it now remains to prove that $\sup_{g\in S_{G}}{\|\alpha(g,.)\|_{\psi}}$ is finite. We use the same techniques as in [2, Appendix A.2 (before Lemma A.1)]. For every $g\in G$, we abusively write